# Bias Is a Subspace, Not a Coordinate: A Geometric Rethinking of Post‑hoc Debiasing in Vision-Language Models

### v1

@@ Abstract
Vision-Language Models (VLMs) have become indispensable for multimodal reasoning, yet their representations often encode and amplify demographic biases, resulting in biased associations and misaligned predictions in downstream tasks. Such behavior undermines fairness and distorts the intended alignment between vision and language. Recent post-hoc approaches attempt to mitigate bias by replacing the most attribute-correlated embedding coordinates with neutral values. However, our systematic analysis reveals three critical failures of this coordinate-wise approach: feature entanglement, poor cross-dataset generalization, and incomplete bias removal. We find that bias is not localized to a few coordinates but is instead distributed across a few linear subspaces. To address these limitations, we propose Subspace Projection Debiasing (SPD), a geometrically principled framework that identifies and removes the entire subspace of linearly decodable bias while reinserting a neutral mean component to preserve semantic fidelity. Extensive experiments across zero-shot classification, text-to-image retrieval, and image generation validate the effectiveness of SPD: our method achieves more robust debiasing with an average improvement of 18.5% across four fairness metrics, while maintaining minimal loss in task performance compared to the best debiasing baseline.
@@ 1 Introduction
Vision-Language Models (VLMs) have rapidly become central to modern multimodal AI, powering image-text retrieval, visual question answering, image captioning, and text-to-image generation [35, 27, 22, 30, 50]. Their broad generalization and emergent cross-modal alignment make them indispensable foundation models [2, 27, 28, 8, 31, 47]. However, an expanding body of work shows that VLMs also inherit and amplify demographic and social biases present in large-scale web data [15, 51, 16, 23]. These biases appear as gendered profession associations [43, 11, 18, 46], racially skewed retrieval rankings [15, 26, 32], or stereotypical captions [51, 6, 36, 52], thereby undermining fairness, reliability, and the trustworthiness of model predictions [53]. Beyond ethical concerns, biased internal representations induce spurious correlations that degrade robustness and cross-domain generalization, as models exploit socially biased features rather than semantically relevant features [44, 12, 38, 45]. Thus, effective debiasing is essential not only for social responsibility but also for preserving the generalization integrity of multimodal systems.
@@
However, our systematic reproduction study (Sec. 3) reveals that these assumptions do not hold in practice. We find that the most representative embedding dimensions for different attributes exhibit substantial overlap. This leads to feature entanglement, as removing one attribute’s dimensions unintentionally distorts representations of others. We show that the indices of important dimensions for specific attributes shift across datasets, undermining SFID’s cross-dataset transferability. Finally, we demonstrate that replacing the top-$m$ most important dimension coordinates leaves measurable residual bias, as attribute information is distributed sparsely across far more dimensions than the imputation process targets [3].
@@ 2 Related works
Training-based Debiasing Methods Prior work on VLM debiasing includes some methods that rely on additional training or fine-tuning to suppress sensitive attribute information from learned representations [4, 39, 19, 29, 40, 1, 21, 9, 33]. Berg et al. [4] optimize learnable prompt tokens through an adversarial objective to decorrelate similarity scores from demographic attributes, while Seth et al. [39] introduce a residual correction module that learns additive residuals to remove bias-predictive components from image embeddings. Although these training-based methods can substantially reduce bias, they require heavy hyperparameter tuning and often fail to transfer reliably across datasets. In addition, most are designed to binary attributes and struggle in handling multi-attribute cases where factors such as race, gender, and age interact.
@@
SFID and Relevant Approaches Among post-hoc approaches, Jung et al. [24] propose SFID, which identifies embedding coordinates most correlated with sensitive attributes and replaces them with neutral statistics estimated from low-confidence samples. This method provides a training-free and computationally efficient solution. However, SFID assumes that bias is localized to specific embedding dimensions and can be mitigated through coordinate-level replacement, which oversimplifies the distributed and entangled structure of multimodal feature spaces. Zhang et al. [49] and Hirota et al. [20] introduce subspace-based neutralization strategies that aim to preserve modality alignment while reducing demographic biases. However, methods such as Zhang et al. [49] rely on attribute supervision, while Hirota et al. [20] are specific to CLIP and thus less transferable to other VLM architectures. Overall, despite the progress of recent post-hoc methods, the field still lacks a unified, training-free framework capable of mitigating bias consistently across modalities while preserving semantic alignment.
@@ 3 Rethinking SFID
We conduct a systematic reproduction of SFID [24] and identify three fundamental assumptions underlying its coordinate-wise design. Through diagnostic experiments, we demonstrate that all three assumptions are violated in practice, motivating our subspace-projection method SPD.
@@ 3.2 Feature Entanglement
To test this, we train three independent Random Forest classifiers on FairFace [25] CLIP [35] ViT-B/32 [10] embeddings (D=512D{=}512), one each for Age, Gender, and Race, and extract the top-m=100m{=}100 dimensions per classifier. We then compute pairwise intersections of these index sets.
@@ A $\cap$ G G $\cap$ R A $\cap$ R A $\cap$ G $\cap$ R
Table 1 reveals severe entanglement among attributes, directly contradicting assumption A1. When SFID replaces 100 dimensions to debias gender, it inevitably disrupts dimensions that are also critical for representing other attributes such as race and age. The results indicate that the same embedding dimensions often carry information about multiple attributes, rather than being cleanly separated. Given that FairFace contains only three labeled attributes, the selected dimensions may additionally encode nuisance factors such as illumination or pose. Consequently, blindly replacing those coordinates as SFID does injects irrelevant noise and simultaneously discards useful semantic information that happens to be co-located in the same dimensions.
@@ 3.3 Cross-Dataset Dimension Drift
SFID’s transferability relies on the assumption that the dimensions identified as most important for a given attribute on one dataset should largely align with those on another dataset. If true, imputation values $\bar{x}_{\text{low}}$ learned from FairFace low-confidence samples can be directly applied to FACET [14] embeddings by replacing the same coordinate indices. We extract the top-$m$ dimensions most important for gender from FairFace and FACET independently, using identical Random Forest training procedures, and measure the intersection size. As a baseline, randomly selecting two $m$-subsets from D=512D{=}512 dimensions yields an expected overlap of $m^{2}/D\approx 4.9$ when m=50m{=}50.
@@
Table 3: Accuracy on FairFace under attribute-replacement (race, gender, or age) settings.“Origin” reports accuracy on the original vanilla embeddings, while the remaining columns report SFID with $m=100$ and SPD with $r=1,5,10$ removed subspace directions.
@@ 3.4 Incomplete Debiasing
SFID assumes that bias is concentrated in the top-$m$ dimensions, such that replacing them should render the attribute unpredictable from the modified embeddings. If this assumption holds, a linear classifier trained on embeddings debiased by SFID should achieve accuracy near the random baseline (50% for binary gender, 14.3% for 7-class race, 11.1% for 9-class age). We apply SFID with m=100m{=}100 to FairFace embeddings for each attribute individually, following the official setup. We then train logistic regression probes on both the original and debiased embeddings to assess information removal effectiveness by comparing their test accuracies, as shown in Tab. 3. Across all attributes, SFID reduces accuracy by less than 1% relative to the original embeddings while remaining far above random baselines, indicating that the target attribute information is still largely preserved in the debiased embeddings. This observation violates assumption A3. Even replacing 100 out of 512 dimensions (19.5% of the embedding) barely affects attribute prediction, suggesting that sensitive information is sparsely distributed across far more than $m$ coordinates. Prior work on neural representations [3] has shown that information is often redundantly encoded; SFID’s top-$m$ selection therefore captures only the most concentrated signal, leaving substantial residual leakage that downstream classifiers can still exploit.
@@ 3.5 Implications for Debiasing Design
Our experiments show that all three assumptions underlying SFID’s coordinate-wise design are systematically violated. These failures stem from a fundamental modeling mismatch: SFID treats bias as coordinate-sparse and dataset-invariant, but our findings reveal that bias is subspace-structured, entangled, and distribution-dependent. Consequently, discrete coordinate replacement cannot achieve robust debiasing: it corrupts entangled dimensions, fails to generalize under distribution shift, and removes only partial bias. A principled solution must instead identify and project onto the null space of all attribute-predictive directions, which motivates our SPD framework in Sec. 4.
@@ 4.1 Overview
Motivated by the three violations identified in Sec. 3, we propose a subspace-projection framework SPD that addresses these limitations. Instead of editing individual coordinates, we model bias as encoded along linear directions in the embedding space and project representations onto their orthogonal complement, removing all linearly predictable attribute information while preserving semantics.
Given a frozen embedding matrix $X=[x_{1},\dots,x_{N}]^{\top}\in\mathbb{R}^{N\times D}$ and sensitive-attribute labels $y\in\{1,\dots,C\}$, our framework consists of three stages: (1) Bias subspace identification, using iterative logistic classifiers to extract directions in the embedding space that are predictive of sensitive attributes; (2) Subspace projection, which removes embedding components along these directions; and (3) Neutral reinjection, where a low-confidence mean (computed via Random Forest, following SFID’s setup) is added back to maintain on-manifold semantics. This procedure is fully post-hoc, requires no retraining, and applies uniformly to any frozen encoder or decoder. Figure 1 illustrates the complete pipeline.
@@ 4.2 Bias Subspace Identification via INLP
where $d_{b}=\sum_{t}\mathrm{rank}\!\left(U^{(t)}\right)$. By construction, no linear classifier can recover the attribute from $X^{(T)}$, ensuring that all linearly decodable bias has been removed. In the simplest case where the iteration terminates after a single step ($T=1$), INLP reduces to learning a single discriminative direction using a linear classifier. However, real-world embeddings typically encode sensitive information across multiple correlated directions, and the multi-step INLP procedure provides a more complete and robust removal.
@@ 4.3 Subspace Projection and Neutral Reinjection
This projection guarantees that $x^{\prime}$ lies entirely within the null space of the attribute classifier, thereby removing all linearly decodable information about the protected attribute. Although orthogonal projection protects linear invariance, it may also remove task-relevant components when attribute and task signals are partially entangled. To stabilize semantics, we add back a neutral baseline along the removed subspace. Following SFID [24], we employ a Random Forest classifier to estimate prediction confidence for each sensitive attribute. For each attribute, we rank samples by confidence and compute the mean embedding of the bottom $\tau$% (low-confidence) instances, denoted as $\bar{x}_{\text{low}}$, which serves as this neutral baseline. The final representation is
@@ 4.4 Discussion and Properties
Our subspace-projection framework offers several desirable properties: (i) training-free and efficient (closed-form operations with lightweight linear probes); (ii) theoretically grounded (INLP provably removes crucial linearly decodable bias directions); (iii) semantically robust (neutral reinjection mitigates overcorrection); (iv) interpretable and controllable (each learned axis is an explicit, human-inspectable direction); and (v) generalizable (the same operation applies to image, text, and multimodal components without modification). In summary, our method transforms discrete coordinate editing into a principled subspace operation that removes attribute-specific information while preserving semantics, providing a unified, interpretable, and efficient post-hoc debiasing solution for VLMs.
@@ 5.1 Evaluating SPD on SFID’s Limitations
To verify how SPD addresses the limitations discussed in Sec. 3, we first evaluate its ability to reduce feature entanglement and achieve more complete debiasing, following the same probe-based setup in Sec. 3.4. Specifically, we apply SPD to FairFace image embeddings and vary the number of removed subspace directions ($r\in\{1,5,10\}$) for each attribute (race, gender, and age). After projection, we train linear probes on the debiased embeddings to assess whether residual attribute information remain.
@@ 5.2 Downstream Tasks
Following Jung et al. [24], we evaluate task performance on three VLM tasks: zero-shot classification, text-to-image retrieval, and text-to-image generation. For each task, we use the same evaluation setup and substitute our debiased encoder for the original model. For our SPD method, the number of removed subspace directions is set to 5 across all downstream tasks.
@@ 5.2.1 Zero-shot Classification
Table 4: Combined experimental results for zero-shot classification (FACET) and text-to-image retrieval (Flickr30K). Bold indicates the best result for each baseline, while underline denotes the second-best result. Improvement (%) is computed against the baseline of the same model: $(\text{baseline}-\text{method})/\text{baseline}\times 100$ for bias metrics $\Delta$ DP and Skew@100 (lower is better).
@@
where $\hat{Y}$ is the predicted class, $c\in\mathcal{C}$ is a demographic class, and $a\in\{0,1\}$ is the sensitive attribute. This captures the average demographic disparity in per-class recall, i.e., how much one group is recognized as class $c$ more often than another. Lower $\Delta DP$ indicates less disparity.

### v2

@@ Abstract
Vision-Language Models (VLMs) have become indispensable for multimodal reasoning, yet their representations often encode and amplify demographic biases, resulting in biased associations and misaligned predictions in downstream tasks. Such behavior undermines fairness and distorts the intended alignment between vision and language. Recent post-hoc approaches attempt to mitigate bias by replacing the most attribute-correlated embedding coordinates with neutral values. However, our systematic analysis reveals three critical limitations of this coordinate-wise approach: feature entanglement, poor cross-dataset generalization, and incomplete bias removal. We find that bias is not localized to a few coordinates but is instead distributed across a few linear subspaces. To address these limitations, we propose Subspace Projection Debiasing (SPD), a geometrically principled framework that identifies and removes the entire subspace of linearly decodable bias while reinserting a neutral mean component to preserve semantic fidelity. Extensive experiments across zero-shot classification, text-to-image retrieval, and image generation validate the effectiveness of SPD: our method achieves more robust debiasing with an average improvement of 18.5% across four fairness metrics, while maintaining minimal loss in task performance compared to the best debiasing baseline.
@@ 1 Introduction
Vision-Language Models (VLMs) have rapidly become central to modern multimodal AI, powering downstream tasks like image-text retrieval, visual question answering, image captioning, text-to-image generation, cross-domain adaptation, multimodal reasoning, and more [38, 29, 23, 56, 32, 52, 55]. Their broad generalization and emergent cross-modal alignment make them indispensable foundation models [2, 29, 28, 8, 34, 51]. However, an expanding body of work shows that VLMs also inherit and amplify demographic and social biases present in large-scale web data [15, 57, 16, 24]. These biases appear as gendered profession associations [48, 11, 17, 19, 50], racially skewed retrieval rankings [15, 27, 35], or stereotypical captions [57, 6, 39, 58], thereby undermining fairness, reliability, and the trustworthiness of model predictions [59]. Beyond ethical concerns, biased internal representations induce spurious correlations that degrade robustness and cross-domain generalization, as models exploit socially biased features rather than semantically relevant features [47, 12, 41, 49, 44]. Thus, effective debiasing is essential not only for social responsibility but also for preserving the generalization integrity of multimodal systems.
@@
However, our systematic reproduction study (Sec. 3) reveals that these assumptions do not hold in practice. We find that the most representative embedding dimensions for different attributes exhibit substantial overlap. This leads to feature entanglement, as removing one attribute’s dimensions unintentionally distorts representations of others. We show that the indices of important dimensions for specific attributes shift across datasets, undermining SFID’s cross-dataset transferability. Finally, we show that replacing the top-$m$ most important coordinates leaves substantial linearly decodable attribute signal, indicating that attribute information is broadly distributed across more dimensions than the imputation process targets [3].
@@ 2 Related works
Training-based Debiasing Methods Prior work on VLM debiasing includes some methods that rely on additional training or fine-tuning to suppress sensitive attribute information from learned representations [4, 42, 21, 30, 43, 1, 22, 9, 36, 31]. Berg et al. [4] optimize learnable prompt tokens through an adversarial objective to decorrelate similarity scores from demographic attributes, while Seth et al. [42] introduce a residual correction module that learns additive residuals to remove bias-predictive components from image embeddings. Although these training-based methods can substantially reduce bias, they require heavy hyperparameter tuning and often fail to transfer reliably across datasets. In addition, most are designed for binary attributes and struggle to handle multi-attribute cases in which factors such as race, gender, and age interact.
@@
SFID and Relevant Approaches Among post-hoc approaches, Jung et al. [25] propose SFID, which identifies embedding coordinates most correlated with sensitive attributes and replaces them with neutral statistics estimated from low-confidence samples. This method provides a training-free and computationally efficient solution. However, SFID assumes that bias is localized to specific embedding dimensions and can be mitigated through coordinate-level replacement, which oversimplifies the distributed and entangled structure of multimodal feature spaces. Zhang et al. [54] and Hirota et al. [20] introduce subspace-based neutralization strategies that aim to preserve modality alignment while reducing demographic biases. However, methods such as Zhang et al. [54] rely on attribute supervision, while Hirota et al. [20] are specific to CLIP and thus less transferable to other VLM architectures. Overall, despite the progress of recent post-hoc methods, the field still lacks a unified, training-free framework capable of mitigating bias consistently across modalities while preserving semantic alignment [33].
@@ 3 Rethinking SFID
We conduct a systematic reproduction of SFID [25] and identify three fundamental assumptions underlying its coordinate-wise design. Through diagnostic experiments, we show that these assumptions do not consistently hold in practice, motivating our subspace-projection method SPD.
@@ 3.2 Feature Entanglement
To test this, we train three independent Random Forest classifiers on FairFace [26] CLIP [38] ViT-B/32 [10] embeddings (D=512D{=}512), one each for Age, Gender, and Race, and extract the top-m=100m{=}100 dimensions per classifier, following the official SFID repository default setting. We then compute pairwise intersections of these index sets.
@@ A $\cap$ G G $\cap$ R A $\cap$ R A $\cap$ G $\cap$ R
Table 1 reveals severe entanglement among attributes, directly contradicting assumption A1. When SFID replaces 100 dimensions to debias gender, it inevitably disrupts dimensions that are also critical for representing other attributes such as race and age. The results indicate that the same embedding dimensions often carry information about multiple attributes, rather than being cleanly separated. Given that FairFace contains only three labeled attributes, the selected dimensions may additionally encode nuisance factors such as illumination or pose. Consequently, blindly replacing those coordinates as SFID does inject irrelevant noise and simultaneously discards useful semantic information that happens to be co-located in the same dimensions.
@@ 3.3 Cross-Dataset Dimension Drift
SFID’s transferability relies on the assumption that the dimensions identified as most important for a given attribute on one dataset should largely align with those on another dataset. If true, imputation values $\bar{x}_{\text{low}}$ learned from FairFace low-confidence samples can be directly applied to FACET [14] embeddings by replacing the same coordinate indices. We extract the top-$m$ dimensions most important for gender from FairFace and FACET independently, using identical Random Forest training procedures, and measure the intersection size. As a baseline, randomly selecting two $m$-subsets from D=512D{=}512 dimensions yields an expected overlap of $m^{2}/D$ ($\approx 4.9$ for m=50m{=}50, $\approx 19.5$ for m=100m{=}100).
@@
Table 3: Linear-probe accuracy for predicting protected attributes (Race/Gender/Age) from FairFace embeddings after applying each intervention. “Origin” reports accuracy on the original vanilla embeddings, while the remaining columns report SFID with $m=100$ and SPD with $r=1,5,10$ removed subspace directions.
@@ 3.4 Incomplete Debiasing
SFID assumes that bias is concentrated in the top-$m$ dimensions, such that replacing them should substantially reduce attribute-related information in the modified embeddings. If this assumption holds, replacing the top-$m$ dimensions should at least eliminate the attribute-related signal accessible to linear models, so that a linear classifier achieves accuracy near the random baseline (50% for binary gender, 14.3% for 7-class race, 11.1% for 9-class age). We apply SFID with m=100m{=}100 to FairFace embeddings for each attribute individually, following the official setup. We then train logistic regression probes on the original and debiased embeddings to assess information removal effectiveness by comparing test accuracies, as shown in Tab. 3. Across all attributes, SFID reduces accuracy by less than 1% relative to the original embeddings while remaining far above random baselines, indicating that a substantial portion of the target attribute remains linearly decodable from the modified representations. Even replacing 100 out of 512 dimensions (19.5% of the embedding) barely affects attribute prediction, suggesting that sensitive information is broadly distributed across far more than $m$ coordinates. Prior work on neural representations [3] has shown that information is often redundantly encoded; SFID’s top-$m$ selection therefore captures only the most concentrated signal. These results suggest that A3 may not strictly hold in practice.
@@ 3.5 Implications for Debiasing Design
Our experiments show that all three assumptions underlying SFID’s coordinate-wise design do not consistently hold in practice. These failures stem from a fundamental modeling mismatch: SFID treats bias as coordinate-sparse and dataset-invariant, but our findings reveal that bias is subspace-structured, entangled, and distribution-dependent. Consequently, discrete coordinate replacement can be brittle under feature entanglement and distribution shift, and may leave residual linearly decodable bias in our experiments. These observations motivate considering subspace-structured interventions. A principled solution must identify and project onto the null space of all attribute-predictive directions, motivating our SPD framework in Sec. 4.
@@ 4.1 Overview
Motivated by the three violations identified in Sec. 3, we propose a subspace-projection framework SPD that addresses these limitations. Instead of editing individual coordinates, we model bias as encoded along linear directions in the embedding space and project representations onto the orthogonal complement of a learned attribute-predictive subspace, suppressing linearly decodable attribute information. Because attributes and semantic factors can be partially entangled, this operation can induce collateral utility loss; the projection depth $r$ is therefore treated as a controllable fairness-utility knob.
Given a frozen embedding matrix $X=[x_{1},\dots,x_{N}]^{\top}\in\mathbb{R}^{N\times D}$ and sensitive-attribute labels $y\in\{1,\dots,C\}$, our framework consists of three stages: (1) Bias subspace identification, which uses iterative logistic classifiers to extract directions in the embedding space that are predictive of sensitive attributes; (2) Subspace projection, which removes embedding components along these directions; and (3) Neutral reinjection, which adds back a low-confidence mean (computed via Random Forest as in SFID) to maintain on-manifold semantics. This procedure is fully post-hoc, requires no retraining, and applies to any frozen encoder or decoder. Figure 1 illustrates the complete pipeline.
@@ 4.2 Bias Subspace Identification via INLP
where $d_{b}=\sum_{t}\mathrm{rank}\!\left(U^{(t)}\right)$. When INLP is run until a linear probe reaches chance (after $T$ iterations), the resulting representation prevents linear attribute recovery beyond chance. In practice, we often remove a limited number of directions $r$ to balance bias suppression and downstream utility. For $T=1$, INLP reduces to learning a single discriminative direction using a linear classifier. However, real-world embeddings encode sensitive information across multiple correlated directions, and the multi-step INLP procedure provides a more complete and robust removal.
@@ 4.3 Subspace Projection and Neutral Reinjection
Projecting onto the orthogonal complement of $U$ removes the component of $x$ in $span(U)$, thereby reducing linear decodability of the protected attribute along the identified directions. When INLP is run to convergence (until a probe reaches chance), this procedure is designed to make the attribute linearly unpredictable; with a finite number of removed directions $r$, residual linear signal may remain and stronger removal can increase collateral semantic loss. Although orthogonal projection protects linear invariance, it may also remove task-relevant components when attribute and task signals are partially entangled. To stabilize semantics, we add back a neutral baseline along the removed subspace. Following SFID [25], we employ a Random Forest classifier to estimate prediction confidence for each sensitive attribute. For each attribute, we rank samples by confidence and compute the mean embedding of the bottom $\tau$% (low-confidence) instances, denoted as $\bar{x}_{\text{low}}$, which serves as this neutral baseline. The final representation is
@@ 4.4 Discussion and Properties
Our subspace-projection framework offers several desirable properties: (i) training-free and efficient (closed-form operations with lightweight linear probes); (ii) theoretically grounded (INLP provably removes crucial linearly decodable bias directions); (iii) semantically robust (neutral reinjection mitigates overcorrection); and (iv) interpretable and controllable (each learned axis is an explicit, human-inspectable direction). In summary, our method transforms discrete coordinate editing into a principled subspace operation that removes attribute-specific information while preserving semantics, providing a unified, interpretable, and efficient post-hoc debiasing solution for VLMs.
@@ 5.1 Evaluating SPD on SFID’s Limitations
To verify how SPD addresses the limitations discussed in Sec. 3, we first evaluate its ability to reduce feature entanglement and achieve more complete debiasing, following the same probe-based setup in Sec. 3.4. Specifically, we apply SPD to FairFace image embeddings and vary the number of removed subspace directions ($r\in\{1,5,10\}$) for each attribute (race, gender, and age). After projection, we train linear probes on the debiased embeddings to assess whether residual attribute information remains.
@@ 5.2 Downstream Tasks
Following Jung et al. [25], we evaluate performance on three VLM tasks: zero-shot classification, text-to-image retrieval, and text-to-image generation. For each task, we use the same evaluation setup and substitute our debiased encoder for the original model. For SPD, the number of removed subspace directions is set to 5 for all tasks.
@@ 5.2.1 Zero-shot Classification
Table 4: Combined experimental results for zero-shot classification (FACET) and text-to-image retrieval (Flickr30K). Bold indicates the best result for each downstream task, while underline denotes the second-best result. Improvement (%) is computed against the baseline of the same model: $(\text{baseline}-\text{method})/\text{baseline}\times 100$ for bias metrics $\Delta$ DP and Skew@100 (lower is better).
@@
where $\hat{Y}$ is the predicted class, $c\in\mathcal{C}$ is a task class, and $a\in\{0,1\}$ is the sensitive attribute. This captures the average demographic disparity in per-class prediction rates, i.e., how much more often one group is predicted as class $c$ than another. Lower $\Delta DP$ indicates less disparity.